HH Co. purchased equipment at the beginning of 2020 for $53,000 with an estimated residual value of $5,300 that it has deprecPart (b) only
Answer #1

Depreciation for 2020 = 53,000*40% = 21,200

Depreciation for 2021 = 31,800*40% = 12,720

Accumulated Depreciation = 21200+12720 = 33,920

Cash 21,200
Accumulated Depreciation 33,920
Gain on disposal 2120
Equipment 53,000